please help!!!!! Flossy ‘Foo Foo’ Fanelli reaches a speed of 350 km/h from rest in 6.2 s in her dragster. If the car is 800 kg and generates a driving force of 16 000 N, find the force of friction
acting on the car.
1
Expert's answer
2020-04-13T10:10:17-0400
"ma=m\frac{\Delta v}{\Delta t}=F-F_f"

"800\frac{350}{6.2\cdot 3.6}=16000-F_f"

"F_f=3500\ N"